A business just borrowed $500,000 to build a new restaurant. The loan terms call for equal annual payments at the end of each year. The loan is for 10 years at an APR of 10.00 percent. How much of the first annual payment will be used to reduce the principal balance?
Answer #1

I have answered the question below

Please up vote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

P = PV*r / [1-(1+r)^-n]

P = 500000*0.1 / [1-(1+0.1)^-10]

P = 50000/ [1-0.3855]

P = 50000/ 0.6144

P = 81372.70

First Repayment = 81372.70

Interest for first year = 500000*0.1 = 50000

Repayment Principal Amount = 31372.70